# The role of the irreducible mass in repetitive Penrose energy extraction processes in a Kerr black hole

### v1

@@ IV The repetitive nonlinear process
We start from a BH with mass $M_{0}$ and angular momentum $L_{0}$ and drop from infinity a particle with energy $E_{0}$ and angular momentum $p_{\phi 0}$ into the BH ergosphere. We fix the values $E_{0}=\mu_{0}$, $\hat{p}_{\phi 1}=-19.434$, $\mu_{2}/\mu_{1}=0.78345$, and we obtain the rest of parameters from Eqs. (10) and (11). The first decay of the sequence has been described in Section III. We can straightforwardly extend Eqs. (12) and (13) to a sequence of decays and evaluate the change of each physical quantity. The BH mass and angular momentum at the beginning of the $n$th process depends on the energy $E_{1,n-1}$ and angular momentum $p_{\phi 1}$ of the particle that the BH captured in the preceding $n-1$ process. Therefore, we have that at the $n$th process
@@
$L_{n}$ $=L_{n-1}+\Delta L_{n-1},\quad\hat{\Delta}L_{n-1}=p_{\phi 1}\mu_{1,n-1}M_{n-1}.$ (16b)
We recall that $p_{\phi 1}$ is a given constant in our calculation strategy, as well as the energy of the incoming particle from infinity, $E_{0}=\mu_{0}=10^{-2}M_{0}$. With the above knowledge, we obtain the changes that occurred in the spin parameter, event horizon, and irreducible mass
@@
The physical explanation of this effect is simple. The irreducible increases its value after each decay and induces an increase in the BH surface area and the horizon radius. When the horizon radius reaches the location where the decays occur, the process stops because, on the horizon, no decay is allowed. Indeed, if one calculates where it exactly occurs by solving the equation $\hat{r}_{+}=\hat{r}$, one obtains again the relation (18). The above behavior is illustrated in Fig. 1 for decays occurring at $\hat{r}=1.2$. The complete set of values of the main parameters for each decay is given in Table 2.
The sequence of decays for the case $\hat{r}=1.5$ reaches saturation after $60$ decays, and $\hat{r}=1.9$ after $205$ decays. Figure 2 shows the final values of the BH properties reached at the saturation point, following the analogous calculation above for decay radii from $\hat{r}=1.2$ to $1.99$. The plot shows that the larger the decay radius, the lower the final spin, so the lower the rotational and extractable energy. Until now, one could think the process is extracting the rotational energy. However, the inspection of the irreducible mass shows it becomes larger and larger, approaching the BH mass and explaining why the extractable energy decreases, approaching zero. Indeed, the extracted energy lowers, also approaching zero. The above demonstrates a dramatic conclusion for the Penrose process: instead of being transferred to the ongoing particle, the BH rotational energy lost mostly converts into irreducible mass.
Figure 2: Final values of the BH properties at the saturation point, for decay radii ranging from $\hat{r}=1.2$ to $1.99$.
@@ V Conclusions
From the above points, we conclude that the repetitive Penrose processes either change very little the BH spin and the rotational energy, leaving most of it to be extracted, or it can approach a final state of a Schwarzschild BH but by converting the BH rotational energy into irreducible mass, without real energy extraction. Therefore, it is mandatory to consider self-consistently for the effect of the irreducible mass increase in the energy extraction processes in the Kerr BH ergosphere.
As a major consequence, we are brought to the quest for alternative BH energy extraction processes that could approach reversibility as much as possible, i.e., keeping the increase of the irreducible mass as small as possible. Electrodynamical processes in the vicinity of a Kerr BH embedded in a magnetic field and surrounded by ionized matter offer a promising alternative (e.g., [5], Rueda and Ruffini, to be submitted).

### v3

@@ IV The repetitive nonlinear process
We start from a BH with mass $M_{0}$ and angular momentum $L_{0}$ and drop from infinity a particle with energy $E_{0}$ and angular momentum $p_{\phi 0}$ into the BH ergosphere. Let us assume the same fixed parameters as above, i.e., $\hat{E}_{0}=1$, $\hat{p}_{\phi 1}=-19.434$, $\nu=0.78345$, and we solve for the other parameters via Eqs. (10). The first decay of the sequence has been described in the previous section. We can straightforwardly extend Eqs. (11) and (12) to a sequence of decays and evaluate the change of each physical quantity. The BH mass and angular momentum at the beginning of the $n$th process depends on the energy $E_{1,n-1}$ and angular momentum $p_{\phi 1}$ of the particle that the BH captured in the preceding $n-1$ process. Therefore, we have that at the $n$th process
@@
$L_{n}$ $=L_{n-1}+\Delta L_{n-1},\quad\Delta L_{n-1}=\hat{p}_{\phi 1}\mu_{1,n-1}M_{n-1}.$ (15b)
With the above knowledge, we obtain the changes in the spin parameter, event horizon, and irreducible mass
@@
The physical explanation of this effect is simple. The irreducible increases its value after each decay and induces an increase in the BH surface area and the horizon radius. When the horizon radius reaches the location where the decays occur, the process stops because, on the horizon, no decay is allowed. Indeed, if one calculates where it exactly occurs by solving the equation $\hat{r}_{+}=\hat{r}$, one obtains again the relation (17). Notice that this result is independent of the initial BH spin. The above behavior is illustrated in Fig. 1 for decays occurring at $\hat{r}=1.2$. The complete set of values of the main parameters for each decay is given in Table 2.
The sequence of decays for the case $\hat{r}=1.5$ reaches saturation after $60$ decays, and $\hat{r}=1.9$ after $205$ decays. Figure 2 shows the final values of the BH properties reached at the saturation point, following the analogous calculation above for decay radii from $\hat{r}=1.2$ to $1.99$. The plot shows that the larger the decay radius, the lower the final spin, so the lower the rotational and extractable energy. Notice that according to Eq. (17), in the limit when the decay radius approaches the ergosphere radius (i.e., $\hat{r}\to 2$), the final spin at the end of the repetitive approaches zero. At this stage, one could think the process is extracting efficiently the rotational energy. However, the inspection of the irreducible mass shows it becomes larger and larger, approaching the BH mass and explaining why the extractable energy decreases, approaching zero. Indeed, the extracted energy lowers, also approaching zero. The above demonstrates a dramatic conclusion for the Penrose process: instead of being transferred to the ongoing particle, the BH rotational energy lost mostly converts into irreducible mass.
Figure 2: Final values of the BH properties when reaches saturation, for decay radii ranging from $\hat{r}=1.2$ to $1.99$.
We have explicitly shown the results for a Kerr BH with initial spin $\hat{a}_{0}=1$. We have verified that given a fixed point of decay, $\hat{r}$, the lower the initial spin, the lower the final extracted energy, the larger the irreducible mass, so the larger the final extractable energy. This result confirms that the example explicitly presented here of an initially maximally rotating Kerr BH is the best condition for energy extraction by repetitive Penrose processes.
@@ V Conclusions
Therefore, the repetitive Penrose processes either change very little the BH spin and the rotational energy, leaving most of it to be extracted (by other means, clearly) or it can approach a final state of a Schwarzschild BH by converting the BH rotational energy into irreducible mass, without real energy extraction.
From all the above, we can conclude that a necessary condition to analyze any energy extraction process from a Kerr BH is the use of the BH mass-energy formula, accounting for the nonlinear relation between the mass and angular momentum, constrained by the irreducible mass value.
Hence, it remains a relevant task to check to which extent the efficiency of the BH energy extraction mechanisms in the literature, e.g., the gravitational Penrose process in its collisional form [10, 11, 12, 13, 14, 15] (see, however, also [16]), or electromagnetic mechanisms based on the Blandford-Znajek mechanism [17], magnetohydrodynamic inflows [18], the electric and magnetic generalizations of the Penrose process (e.g., [19, 20, 21, 22, 23, 24]), or magnetic-reconnection mechanisms [25], are affected by this result.
As a major consequence, we are brought to the quest for alternative BH energy extraction processes that could approach reversibility as much as possible, i.e., keeping the increase of the irreducible mass as small as possible. Electrodynamical processes in the vicinity of a Kerr BH embedded in a magnetic field and surrounded by ionized matter offer a promising alternative (e.g., [26], Rueda and Ruffini, submitted).